OpenStudy (pokemon23):

separate 144 into two parts such hat one part will be 12 less than twice the other. separate 144 into two parts such hat one part will be 12 less than twice the other. @Mathematics

OpenStudy (mertsj):

The two parts are x and 144-x x +12 = 2 (144-x) x+12 = 288 - 2x 3x = 276 x = 92 144-x = 52

OpenStudy (anonymous):

92 and 52.

OpenStudy (lilg132):

x + y = 144 x = 2y - 12 144 = 2y - 12 + y 156 = 2y + y 156 = 3y y = 52 x = 2(52) - 12 x = 92